OpenStudy (anonymous):

Factor: x^2+11xy+28y^2

OpenStudy (anonymous):

7+4=11 7*4 =28 hint

OpenStudy (anonymous):

confused with the xy

OpenStudy (anonymous):

(x 7y ) (x 4y )

OpenStudy (anonymous):

what signs go in there

OpenStudy (anonymous):

+

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

I see that it is (x+7y)(x+4y) i just get confused with 2 variables.

OpenStudy (sriram):

another method if u get confused by the existance of both x & y is that u divided the whole stuff by y^2 then u get a quadratic in x/y u can replace x/y = z n solve quadratic then in solutn write z again as x/y